Consider a project that requires an immediate cash outflow of $100,000 and provides a perpetual annual inflow of $15,000 starting two years from today. The cost of capital is 12 percent. What is the project's PI? 

1.04 

1.12 

1.25 

1.33
Answer #1

Present value at year 1 = Perpetual cash flow / cost of capital

Present value at year 1 = 15,000 / 0.12

Present value at year 1 = 125,000

Present value = 125,000 / (1 + 0.12)

Present value = 111,607.1429

Profitability index = Present value / initila investment

Profitability index = 111,607.1429 / 100,000

Profitability index = 1.12